### v2

# MoSECroT: Model Stitching with Static Word Embeddings for Crosslingual Zero-shot Transfer

Abstract
Transformer-based pre-trained language models (PLMs) have achieved remarkable performance in various natural language processing (NLP) tasks. However, pre-training such models can take considerable resources that are almost only available to high-resource languages. On the contrary, static word embeddings are easier to train in terms of computing resources and the amount of data required. In this paper, we introduce MoSECroT (Model Stitching with Static Word Embeddings for Crosslingual Zero-shot Transfer), a novel and challenging task that is especially relevant to low-resource languages for which static word embeddings are available. To tackle the task, we present the first framework that leverages relative representations to construct a common space for the embeddings of a source language PLM and the static word embeddings of a target language. In this way, we can train the PLM on source-language training data and perform zero-shot transfer to the target language by simply swapping the embedding layer. However, through extensive experiments on two classification datasets, we show that although our proposed framework is competitive with weak baselines when addressing MoSECroT, it fails to achieve competitive results compared with some strong baselines. In this paper, we attempt to explain this negative result and provide several thoughts on possible improvement.
MoSECroT: Model Stitching with Static Word Embeddings
for Crosslingual Zero-shot Transfer
**footnotetext: Equal contribution.
1 Introduction
The emergence of PLMs and their multilingual counterparts (mPLMs) (Devlin et al., 2019; Conneau et al., 2020) have proven effective for various NLP tasks (Artetxe et al., 2020; ImaniGooghari et al., 2023). However, such models are mostly limited to no more than a hundred languages, as the pre-training requires considerable data that is only available to these languages, leaving the majority of the world’s low-resource languages uncovered. In this work, we explore the possibility of leveraging (1) a PLM in a source language, (2) static word embeddings in a target language, which are readily available for many low-resource languages and are much easier to train, and (3) a technique called model stitching, to enable zero-shot on the target language without the need to pre-train.
Our contribution is summarized as follows: (i) we introduce MoSECroT, a novel and challenging task for (especially low-resource) languages where static word embeddings are available. (ii) We propose a solution that leverages relative representations to construct a common space for source (English in our case) and target languages and that allows zero-shot transfer for the target languages.
2 Related Work
Aligned crosslingual word embeddings enable transfer learning by benefiting from a shared representation space for the source and target languages. Such embedding pairs are typically either trained jointly (Hermann and Blunsom, 2014; Vulic and Moens, 2016) or obtained through post-alignment (Lample et al., 2018; Artetxe et al., 2018). Our work applies a transformation in the manner of the latter to align two embedding spaces where the source embeddings are derived from a PLM and target embeddings are static word embeddings.
Based on a recent consensus that similar inner representations are learned by neural networks regardless of their architecture or domain (Kornblith et al., 2019; Vulić et al., 2020), Moschella et al. (2023) propose an approach to align latent spaces with respect to a set of samples, called parallel anchors. They transform the original, absolute space to one defined by relative coordinates of the parallel anchors, and denote all the transformed samples in the relative coordinates as relative representations.
Model stitching was proposed as a way to combine (stitch together) components of different neural models. Trainable stitching layers are first introduced by Lenc and Vedaldi (2015), with a series of subsequent works demonstrating the effectiveness of the approach (Bianchi et al., 2020; Bansal et al., 2021).
3 MoSECroT Task Setting
The task setting is straightforward: given a PLM of a high-resource language (regarded as the source language) and static word embeddings of another language (low-resource and regarded as the target language), the goal is to achieve zero-shot transfer by using the target language embeddings directly with the source language model via embedding layer stitching. This can be done by first applying an alignment between the source and target embedding spaces and subsequently swapping the embedding matrices of the PLM.
We propose a novel method that leverages relative representations for embedding space mapping. In the following, we describe our methodology in more detail.
4 Methodology
Parallel anchor selection
We first extract bilingual parallel lexica between the source and the target language. For most high-resource languages, large bilingual lexica are available from MUSE [^1] https://github.com/facebookresearch/MUSE. For low-resource languages, we crawl translations of source language vocabulary from PanLex [^2] https://panlex.org and Google Translate [^3] https://translate.google.com. Then we derive a subset of the lexica as the parallel anchors $A$ for our method: we only keep those parallel lexica which exist in the embeddings of source and target languages [^4] The source language is always English and its embeddings are extracted from English BERT’s (Devlin et al., 2019) token embeddings. For target languages, embeddings are static word embeddings from fastText (Bojanowski et al., 2017)..
Relative representations
Following Moschella et al. (2023), we build relative representations (RRs) for each token in the embedding space based on their similarities with anchor tokens in the respective language. Specifically, we compute the cosine similarity of the embedding of each token with the embedding of each anchor token. This computation is done in the embedding spaces of the source and target languages respectively. For example, in the source language, the similarity between token $x_{i}$ and anchor $a_{j}$ is calculated as follows:
$r^{s}_{(i,j)}=\text{cos-sim}(\boldsymbol{E}^{s}_{\{x_{i}\}},\boldsymbol{E}^{s}_{\{a_{j}\}})$
where $\boldsymbol{E}^{s}_{\{x_{i}\}}$, $\boldsymbol{E}^{s}_{\{a_{j}\}}$ are the word embedding of $x_{i}$ and $a_{j}$ in the source PLM embeddings $\boldsymbol{E}^{s}$. The relative representation of token $x_{i}$ from the source language is then defined as follows:
$\boldsymbol{R}^{s}_{\{x_{i}\}}=[r^{s}_{(i,1)},r^{s}_{(i,2)},r^{s}_{(i,3)},\cdots,r^{s}_{(i,|A|)}]$
Note that the relative representation is sensitive to the order of the anchors, so the relative representation for each token is computed with the anchors in the same order. This computation results in a matrix $\boldsymbol{R}^{s}\in\mathbb{R}^{|V^{s}|\times|A|}$ of source language embeddings and a matrix $\boldsymbol{R}^{t}\in\mathbb{R}^{|V^{t}|\times|A|}$ of target language embeddings, where $|V^{s}|$ (resp. $|V^{t}|$) is the source-language (resp. target-language) vocabulary size and $|A|$ is the number of parallel anchors.
Embedding mapping
The obtained relative representations are vectors in $\mathbb{R}^{|A|}$ for both source and target languages. This dimension does not suit the hidden dimension of the Transformer body of the source PLM. Therefore, we propose to map the relative representations of both source and target languages back to $\mathbb{R}^{D}$, which is the same as the dimension of $\boldsymbol{E}^{s}$. Given $\boldsymbol{E}^{s}$ and $\boldsymbol{R}^{s}$ for source language (resp. $\boldsymbol{E}^{t}$ and $\boldsymbol{R}^{t}$ for target language), we compute the transformed embedding of any token $x_{i}$ from the source language (resp. any token $y_{i}$ from the target language) as follows:
$\boldsymbol{F}^{s}_{\{x_{i}\}}=\frac{\sum_{n\in\mathbb{N}(x_{i})}(\boldsymbol{R}^{s}_{\{x_{i}\},n}/\tau\cdot\boldsymbol{E}^{s}_{\{n\}})}{\sum_{n\in\mathbb{N}(x_{i})}\boldsymbol{R}^{s}_{\{x_{i}\},n}/\tau}$
$\boldsymbol{F}^{t}_{\{y_{i}\}}=\frac{\sum_{n\in\mathbb{N}(y_{i})}(\boldsymbol{R}^{t}_{\{y_{i}\},n}/\tau\cdot\boldsymbol{E}^{s}_{\{n\}})}{\sum_{n\in\mathbb{N}(y_{i})}\boldsymbol{R}^{t}_{\{y_{i}\},n}/\tau}$
where $\mathbb{N}(x_{i})$ (resp. $\mathbb{N}(y_{i})$) is the set of top-$k$ closest anchors in terms of the cosine similarity recorded in $\boldsymbol{R}^{s}_{x_{i}}$ (resp. $\boldsymbol{R}^{t}_{y_{i}}$), $\boldsymbol{R}^{s}_{\{x_{i}\},n}$ (resp. $\boldsymbol{R}^{s}_{\{y_{i}\},n}$) is the cosine similarity between $\boldsymbol{E}^{s}_{\{x_{i}\}}$ (resp. $\boldsymbol{E}^{t}_{\{y_{i}\}}$) and $\boldsymbol{E}^{s}_{\{n\}}$ (resp. $\boldsymbol{E}^{t}_{\{n\}}$), and $\tau$ is the temperature. Note that both the resulting transformed embeddings $\boldsymbol{F}^{s}_{\{x_{i}\}}$ and $\boldsymbol{F}^{t}_{\{y_{i}\}}$ are in $\mathbb{R}^{D}$, because it is a weighted sum of the anchor embedding in the source language, i.e., $\boldsymbol{E}^{s}_{\{n\}}$. A simple summary of the process is to represent any token, no matter whether it is from the source or target language, as a weighted sum of the embeddings of some parallel anchors in the source-language embedding space.
Zero-shot stitching
So far we project the target-language embeddings to $\mathbb{R}^{D}$, which suits the hidden dimension of the Transformer body of the source language. We also manipulate the original token embedding matrix of the source language, where the matrix dimensions stay the same: $\boldsymbol{F}^{s}\in\mathbb{R}^{|V^{s}|\times D}$. We can simply fine-tune the model ($\boldsymbol{F}^{s}$ and the Transformer body) on the source-language train set of a downstream task and then assemble a target-language model for zero-shot transfer, without training on the target language. To do this, we only need to swap the source-language embeddings $\boldsymbol{F}^{s}$ with target-language embeddings $\boldsymbol{F}^{t}$.
5 Experiments
5.1 Setup
We use the cased version of the English BERT model (bert-base-cased) as the source language PLM and consider eight target languages. Three of the target languages are high-resource: German (de), Spanish (es), and Chinese (zh), and the rest are low-resource: Faroese (fo), Maltese (mt), Eastern Low German (nds), Sakha (sah), and Tatar (tt). Pre-trained static embeddings for all target languages are available from fastText [^5] https://fasttext.cc/docs/en/pretrained-vectors.html, except for Eastern Low German, for which we download fastText embeddings from Huggingface [^6] https://huggingface.co/facebook/fasttext-nds-vectors.
Using the method proposed in §4, we obtain pairwise parallel anchors between English and each target language. The size of the anchor set varies depending on the vocabulary size of the language’s embeddings and the overlap between the English and target language lexica, which is the following for each target language: 11836 (en-de), 11395 (en-es), 7662 (en-zh), 1577 (en-fo), 2600 (en-mt), 1309 (nds), 3242 (en-sah), and 9275 (en-tt).
We evaluate the proposed method on two text classification datasets: Multilingual Amazon Reviews Corpus (Keung et al., 2020) and Taxi1500 (Ma et al., 2023). See §C for details.
Apart from the standard weighting scheme illustrated in §4, we propose two more settings: one where we apply softmax over relative representation weights (in the Embedding mapping step), and another using sparsemax (Martins and Astudillo, 2016). Compared to softmax, sparsemax produces sparse weight distributions, meaning more similarities are concentrated on fewer anchors. We conduct preliminary experiments to identify the optimal top-$k$ closest anchors $\in\{1,10,50,100\}$ and find that the results are best when using the top 50 anchors. See §A for an exploration of how different choices of $k$ influence the performance.
de es zh
LR 0.52 0.51 0.50
mBERT 0.61 0.65 0.51
LS 0.46 0.46 0.30
RRs standard top-50 0.53 0.51 0.38
RRs softmax top-50 0.50 0.53 0.38
RRs sparsemax top-50 0.56 0.57 0.24
Table 1: Evaluation results on the Amazon Reviews Corpus. We report macro $F_{1}$ scores on the test sets of three high-resource target languages. Bold: highest score per column.
de es zh mt sah fo nds tt
LR 0.30 0.32 0.56 0.38 0.48 0.47 0.18 0.43
mBERT 0.24 0.60 0.62 0.08 0.07 0.18 0.12 0.18
LS 0.14 0.26 0.24 0.08 0.12 0.06 0.08 0.07
RRs standard top50 0.20 0.44 0.28 0.14 0.16 0.16 0.06 0.14
RRs softmax top50 0.20 0.48 0.28 0.15 0.19 0.16 0.06 0.17
RRs sparsemax top50 0.24 0.37 0.13 0.15 0.18 0.20 0.13 0.21
Table 2: Evaluation results on the Taxi1500 dataset. Reported metrics are macro $F_{1}$ scores on the test sets of eight target languages. Scores are averaged over five runs with different random seeds. Bold: highest score per column.
5.2 Baselines
We compare our method against three baselines:
Logistic Regression (LR)
We train a simple target language logistic regression classifier using the average of static word embeddings of the input sentences. This approach does not require expensive training of a language model but assumes we have sufficient target language training data for a specific downstream task, which is hardly the case for most low-resource languages in real scenarios.
mBERT
We fine-tune multilingual BERT (mBERT) (Devlin et al., 2019), which is pre-trained on more than 100 languages, using the English training data, and perform zero-shot predictions directly on the target language test data.
Least squares projection (LS)
We propose a straightforward approach, inspired by embedding alignment frameworks such as VecMap (Artetxe et al., 2018), to project target language embeddings into the same space as the English PLM embeddings. Specifically, we learn a transformation matrix $\boldsymbol{W}\in\mathbb{R}^{D^{t}\times D}$ by minimizing $||\boldsymbol{A}^{t}\boldsymbol{W}-\boldsymbol{A}^{s}||_{F}^{2}$, where $\boldsymbol{A}^{t}\in\mathbb{R}^{|A|\times D^{t}}$ is the embeddings of anchors in the target language and $\boldsymbol{A}^{s}\in\mathbb{R}^{|A|\times D}$ is the embeddings of anchors from the English PLM. We then project all target language embeddings using $\boldsymbol{W}$ and replace the BERT embedding layer with the resulting matrix.
5.3 Results
We present evaluation results of RRs with the proposed settings (§5.1) and compare them with the baselines in Tables 1 and 2. Macro $F_{1}$ is used due to class imbalance in both datasets.
We notice that the naive LS baseline is almost always beaten by the proposed method under multiple RR settings on both datasets. The only exception is nds, in Table 2, where both LS and RRs perform badly. This observation is a strong indicator that RRs can better leverage the semantic similarity encoded in different types of embeddings than LS.
Not very surprisingly, zero-shot with mBERT is effective for high-resource languages in both datasets but underperforms LR with large gaps on low-resource languages in Taxi1500. There are two possible explanations for this phenomenon. First, representations in mBERT are not well-aligned across low-resource languages. This is possibly due to data sparsity, which is observed by previous work (Wu and Dredze, 2020), where mBERT archives good performance on high-resource languages but sub-optimal performance on low-resource languages. Second, Taxi1500 is a relatively easy task: a model with good alignment across languages, especially on the word level, is expected to perform well. This argument is supported by a previous work (Liu et al., 2023), where well-aligned word embeddings achieve better zero-shot crosslingual performance than mPLMs on a wide range of languages in Taxi1500.
Although none of the RR settings outperforms mBERT on high-resource languages (as mentioned earlier, mBERT has strong crosslingual transfer ability on high-resource languages), for all five low-resource languages not seen by mBERT (mt, sah, fo, nds, tt), RRs outperform mBERT consistently, with varying margins (ranging from +0.12 for sah to +0.01 for nds). This suggests that RRs can be a promising alternative when a low-resource language is not covered by an mPLM.
6 Analysis
In this section, we want to propose possible reasons for the suboptimal results obtained by our framework tackling the MoSECroT task.
Anchor selection
The quality of the parallel anchors largely relies on the quality of the bilingual lexica, which may contain, among others, polysemous words, that may influence the alignment quality. Normalization can also be a source of ambiguity. For example, MUSE converts all words into lowercase, so the word sie can have three meanings in the German-English lexicon: you, she, and they. We (1) only consider one translation (if there are multiple) for each target language word, which may not be the most accurate one; and (2) treat all target language words whose translations are in the source language vocabulary as anchors, which increases the frequency of noisy translation pairs.
We try to decrease the influence of potentially noisy anchor pairs by reducing the number of anchors to 3000 and 500 (the original anchor set used during the preliminary experiments contains 6731 anchors, see §4) through random sampling, following the observation by Moschella et al. (2023) that uniform selection from an anchor set is both straightforward and has good performance. We also remove stop words, whose translations are more unstable, from the anchor set. Neither of the two modifications shows an improvement over the full anchor set (see §B for the comparison). One possible explanation is that the translation qualities vary across anchors and thus we cannot predict the quality of sampled anchors.
Translation quality
We find that a large portion of translations retrieved from PanLex are of low quality. This is partly due to PanLex using intermediate languages when direct translation is unavailable for the language pair. We filter the translations by empirically setting a threshold to the translation quality scores, available through the API for every translation. Nevertheless, we note that a high translation quality score does not guarantee the translation is perfect, and many translations are good despite having low translation quality scores. We believe the lack of high-quality parallel lexica is a possible reason that RRs do not reach their full potential on low-resource languages.
Reinitialized embedding space
Our method requires swapping the original PLM embeddings with the transformed English RRs before fine-tuning on English data, whereas the embedding space of RRs might diverge substantially from the original embedding space. As a result, it is unclear whether the rest of the model parameters can be adapted to the new embeddings during fine-tuning, especially on smaller datasets like Taxi1500. We thus suggest the alteration of the embedding space through reinitialization with RRs as a likely factor as to why we do not achieve good performance.
7 Conclusion
In this work, we introduce MoSECroT, a novel and challenging task that is relevant for, in particular, low-resource languages for which static word embeddings are available but few resources exist. In addition, we propose for the first time a method that leverages relative representations for embedding space mapping and enables zero-shot transfer. Specifically, we fine-tune a monolingual English language model using only English data, swap the embeddings with target language embeddings aligned using RRs, and apply zero-shot evaluation on the target language. We show that the proposed method is promising compared with mBERT on unseen languages but only modest improvements are achieved. We provide several possible reasons and leave improvement possibilities for future research.
Limitations
In this work, we propose the task of MoSECroT and a solution to leverage available static pre-trained embeddings and tackle downstream tasks for low-resource languages. Our work has a few limitations open to future research. First, we only experiment with one model architecture (BERT). Although many language-specific BERT models exist and thus our method is applicable to a wide range of high-resource source languages, it would nevertheless be interesting to compare performance across different model architectures. Second, the explored tasks are exclusively text classification tasks. We expect that the robustness of our method can be much better studied by applying it to a more diverse set of tasks.
Acknowledgements
This work was funded by the European Research Council (grant #740516).
Appendix A Number of closest anchors
In addition to using all (6731) parallel anchors, we consider only the top-$k$ ($k\in\{1,10,50,100\}$) closest anchors of each word. We identify the optimal value for $k$ closest anchors based on zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (§C.1). Table 3 shows results for different $k$ values.
$k$ de zh
1 0.44 0.41
10 0.51 0.38
50 0.50 0.40
100 0.51 0.38
6731 0.44 0.21
Table 3: Number of closest parallel anchors ($k$) and the corresponding zero-shot performance on de and zh portions of the Amazon Reviews Corpus.
Appendix B Total number of anchors
Following Moschella et al. (2023), we randomly sample a subset of the parallel anchors ($|\boldsymbol{A}|\in\{500,3000\}$), and exclude stop words from the anchor set. Table 4 shows zero-shot performance on German and Chinese portions of the Amazon Reviews Corpus (§C.1).
$|\boldsymbol{A}|$ de zh
500 0.39 0.19
3000 0.19 0.19
6731 0.44 0.21
Table 4: The total number of parallel anchors and the corresponding zero-shot performance on de and zh portions of the Amazon Reviews Corpus.
Appendix C Evaluation datasets
C.1 Multilingual Amazon Reviews Corpus
Presented by Keung et al. (2020) and containing product reviews in six languages, the original dataset uses five labels corresponding to star ratings, which we aggregate into three classes: positive, neutral, and negative. We evaluate the three high-resource target languages (de, es, zh) on this dataset.
C.2 Taxi1500
Taxi1500 (Ma et al., 2023) is a classification dataset containing six classes for more than 1500 languages, including all of our target languages. We follow the authors’ original training procedure and hyperparameters and use a learning rate of 1e-5 instead of 2e-5, which we find works better for our settings.
Appendix D Computational resources
Training can be completed in under three hours on eight NVIDIA GeForce GTX 1080 Ti GPUs for the Multilingual Amazon Reviews Corpus or about half an hour on a single NVIDIA GeForce GTX 1080 Ti GPU for Taxi1500.